vehicles pass through a junction on a busy road at an average rate of 5 per minute.What is the probability that no car will pass in a given minute
1
Expert's answer
2020-03-03T14:29:17-0500

The passage time is a Poisson process with mean 5.

"P(0)=\frac{e^{-5}\times{5^0}}{0!}=e^{-5}=0.006737947"

For this purpose the Excel function POISSON.DIST can be used. The formula is given by =POISSON.DIST(0,5,FALSE) which yields 0.006737947.